use the method of elimination to solve each system of linear equations:

.1x-.4y=-2
.7x- y=-9.6

Answer by checkley77(12844)

.1x-.4y=-2 multiply by -7 & add.
.7x- y=-9.6
-.7x+2.8y=14
-----------------
1.8y=4.4
y=4.4/1.8
y=2.444 ans.
.1x-.4*2.444=-2
.1x-.977=-2
.1x=-2+.977
.1x=-1.023
x=-1.023/.1
x=-10.23 ans.
Proof:
.7*-10.23-2.444=-9.6
-7.161-2.444=-9.6
9.6~9.6